OpenStudy (anonymous):

Using the correct Order of Operations, simplify as much as possible: 3(5-7)^2-6(3)

OpenStudy (anonymous):

3(5-7)^2 - 6(3) = 3(2)^2 - 6(3) = 3(4) - 6(3) = 12 - 18 = -6